cherry:

if the scale factor between two circles is 2x/5y, what is the ratio of their areas?

jhonyy9:

you need to know in reality what mean this scale factor of 2x/5y between these two circle

Mercury:

in general, the ratio of two areas is the square of the linear ratio quick derivation for a circle: let a = pi * r^2 be the smaller circle, and A = pi * R^2 be the larger circle then A/a = (pi * R^2) / (pi * r^2) = R^2/r^2 = (R/r)^2